DESCRIPTION OF THE F ANT ASIO INSTRUMENTAL SET-UP

K. DIDRICHE, P. MACKO and \underline{M. HERMAN}, Laboratoire de Chimie quantique et Photophysique, CP160/09, Université Libre de Bruxelles (U.L.B.), Ave. F.-D. Roosevelt 50, Brussels, B-1050, Belgium.

We have built a new set-up, named F ANT ASIO, for ”Fourier trANsform, Tunable diode and quadrupole mAss spectrometers interfaced to a Supersonic expansIOn”. Initial results will be reported, including the mapping of the jet expansion and the observation of acetylene clusters in the infrared spectral range.